OpenStudy (anonymous):

how would you solve the equation (x+3)^{3/2}-11=53? what is the correct solution?

OpenStudy (anonymous):

\[\begin{align*} (x+3)^{3/2}-11&=53\\ (x+3)^{3/2}&=64\\ \left((x+3)^{3/2}\right)^{2/3}&=(64)^{2/3}\\ x+3&=(64)^{2/3}\\ &~\vdots \end{align*}\]

OpenStudy (anonymous):

\[(x+3)^{3/2}-11=53 \rightarrow (x+3)^{3/2}=53+11 \rightarrow (x+3)^{3/2}=64\] i.e. \[(x+3)=64^{2/3} \rightarrow (x+3)=(4^3)^{2/3} \] i.e. \[ (x+3)=(4^2) \rightarrow x+3=16 \rightarrow x= 16-3 \rightarrow x=13\]